# Quantifying Emergence in Large Language Models

### v1

@@ Abstract
Emergence, broadly conceptualized as the “intelligent” behaviors of LLMs, has recently been studied and proved challenging to quantify due to the lack of a measurable definition. Most commonly, it has been estimated statistically through model performances across extensive datasets and tasks, which consumes significant resources. In addition, such estimation is difficult to interpret and may not accurately reflect the models’ intrinsic emergence. In this work, we propose a quantifiable solution for estimating emergence. Inspired by emergentism in dynamics, we quantify the strength of emergence by comparing the entropy reduction of the macroscopic (semantic) level with that of the microscopic (token) level, both of which are derived from the representations within the transformer block. Using a low-cost estimator, our quantification method demonstrates consistent behaviors across a suite of LMs (GPT-2, GEMMA, etc.) under both in-context learning (ICL) and natural sentences. Empirical results show that (1) our method gives consistent measurements which align with existing observations based on performance metrics, validating the effectiveness of our emergence quantification; (2) our proposed metric uncovers novel emergence patterns such as the correlations between the variance of our metric and the number of “shots” in ICL, which further suggests a new way of interpreting hallucinations in LLMs; (3) we offer a potential solution towards estimating the emergence of larger and closed-resource LMs via smaller LMs like GPT-2. Our codes are available at: https://github.com/Zodiark-ch/Emergence-of-LLMs/.
@@ 1 Introduction
Emergence, one of the most enigmatic and elusive attributes of LLMs [31, 24, 26], is broadly conceptualized as a phenomenon of “high-level intelligence” that becomes observable only when the model reaches a sufficiently large size. It encompasses the comprehension of abstract concepts [33], complex language skills [2], natural human interactions [18], instructions following [36], generalization out of domains [12], and even unanticipated patterns like hallucinations [22].
To measure emergence, existing research adopts a compromised approximation: evaluating the performance of a model across extensive datasets and tasks under a pre-defined performance metric such as accuracy, F1-score, etc. For example, recent work [24], under the linear and continuous metrics, has demonstrated that emergence is not “unreasonable like magic” but rather can be extrapolated from smaller models. However, these evaluations pose three main limitations. Firstly, evaluating emergence through performance metrics across numerous tasks and datasets is time-consuming and resource-intensive. Secondly, it does not provide an internal explanation at a finer-grained level such as specific tokens or model components to better understand the mechanism. Thirdly, pre-defined performance metrics may not correlate precisely with the strength of emergence, especially when the metric fails to capture the creativity or flexibility of LM-generated answers.
Despite these limitations, research into the intrinsic nature of emergence still remains a highly under-explored area. Therefore, our work is dedicated to proposing a quantifiable definition of emergence in LLMs and a low-cost implementation. Specifically, motivated by theories in emergentism [23, 13, 14], we model LLMs as a Markov process and define the emergence strength between internal transformer layers from the perspective of information theory. Concisely, emergence can be described as a process where the entropy reduction of the entire sequence exceeds the entropy reduction of individual tokens. To compute entropy reduction, we resort to mutual information between successive transformer layers and adopt a practically effective estimation algorithm proposed in [6] which is suitable for high-dimensional continuous representations.
To demonstrate the effectiveness and validity of our proposed emergence metric, we devise a suite of comprehensive experiments encompassing two different scenarios. In the first scenario, we target on emergence in “shots” and curate a group of synthetic datasets under the in-context learning (ICL) setting to account for different numbers of demonstrations (“shots”) in ICL as well as different context domains. In the second scenario, we focus on emergence in sentences and collect two benchmark datasets consisting of real-world natural language questions/answers. Under both scenarios, we experiment with different LMs including GPT-2 [21], GEMMA [28], and OpenLlama [10] to study the emergence patterns. Interestingly, we observe robust phenomenon across these LMs when quantifying their emergence, which align with existing findings based on commonly used performance metrics. For instance, an increase in model size and context length correlates with higher emergence. Additionally, novel patterns are detected, e.g., sequences with ICL-style “shots” display distinct emergence patterns compared to natural sequences.
On top of that, our methodology offers underlying insights applicable to several research areas within LLMs: 1) By estimating the emergence strength across numerous LLMs, empirical formulas about emergence akin to Scaling Laws [3] can be developed. 2) Our research indicates a potential connection between the high variations of emergence strength and occurrence of hallucinations in LLMs, paving a future direction towards hallucination detection via models’ internal mechanisms. 3) Our method demonstrates the feasibility of using smaller LMs like GPT-2 to estimate the emergence capabilities of extremely large and close-resource LMs like GPT-4 [1] with low computational cost. Overall, our contributions could be summarized below:
  • To our best knowledge, this is the first work to theoretically quantify the emergence strength in LLMs and provide a rather low-cost solution to estimate the emergence strength.
  • Extensive experiments under the settings of ICL and natural sentences have uncovered numerous inspiring emergence patterns, aligning with and beyond existing findings.
  • Our experimental findings further potentially support the feasibility and encourage the exploration of a wide range of applications of greater value within the scope of LLMs.
@@ 2 Related Work
Initially, the emergence capabilities of LLMs were described as “Emergence is when quantitative changes in a system result in qualitative changes in behavior” [31]. This description indicates that such capabilities are abrupt and unpredictable — eluding extrapolation from smaller-scale models. Therefore, a prevalent approach involves adopting extensive datasets and tasks, then evaluating the models’ performance via metrics—such as accuracy, F1 score, and Exact Match (EM)—which depend on the specific datasets and tasks, and using these metrics to statistically elucidate the emergence capabilities. To date, numerous studies [26, 18, 35, 17, 7, 9] have continued this approach and have observed such an abrupt phenomenon.
More recently, several studies have demonstrated both theoretically and experimentally that emergence capabilities can be observed as smooth and continuous phenomenon. [2] integrated text pieces with language skills into a bipartite random graph thus employing Scaling Laws to model emergence properties ranging from smaller to larger models. [24] proposed that the abrupt nature of emergence derives from nonlinear or discrete metrics, and experimentally confirmed that when metrics are linear and continuous, the emergence capabilities appear smooth and predictable.
Despite a novel pathway for defining and quantifying the emergence in LLMs pioneered by us, our empirical investigations substantiate these previous viewpoints: strength of emergence does scale continuously with model size, datasets, prompt form, and context.
@@ 3 Method
3.1 How to Model Emergence in LLMs
In this paper [^1] Please note that this paper addresses decoder-only autoregressive language models, while other language models that do not meet the requirements, such as BERT [11] family, are not considered within the scope., we identify the transformer block as the fundamental unit for observations of robust emergence phenomenon. Specifically, we employ $l=0,1,\dots,L-1$ to index transformer blocks within a language model, where $L$ represents the total number of blocks. For instance, GPT-2 XL (1.6B parameters)  [^2] https://huggingface.co/openai-community/gpt2-xl comprises 12 blocks ($L=12$), and Gemma-2B  [^3] https://huggingface.co/google/gemma-2b totals 18 blocks ($L=18$). Without loss of generality, we hypothesize that the process of token representations passing through transformer blocks constitutes a Markov process. Specifically, For any transformer block $l$, given an input sequence of token length $T$ and hidden state dimension $D$, the input representation is given by $H_{l}=\{h_{l}^{0},h_{l}^{1},h_{l}^{2},\dots,h_{l}^{T-1}\}$ and the output representation is $H_{l+1}=\{h_{l+1}^{0},h_{l+1}^{1},h_{l+1}^{2},\dots,h_{l+1}^{T-1}\}$, where $H\in\mathbb{R}^{T\times D}$ and $h^{t}\in\mathbb{R}^{1\times D}$.
Figure 1: The analogy of NTP to Markov process. Taking the output representation of token2222 in Block 00 ($h^{2}_{1}$) as an example, which receives information from input representations of $h^{0}_{0}$, $h^{1}_{0}$, and $h^{2}_{0}$,satisfying $p_{h^{2}_{l+1}|H_{l}^{\leqslant t}}=p_{h^{2}_{l+1}|h^{0}_{l},h^{1}_{l},h^{2}_{l}}$.
@@ 3.1 How to Model Emergence in LLMs
The Next Token Prediction (NTP) mechanism undergoes a Markovian stochastic process following a transition probability of any $h^{t}_{l+1}$ with $p_{h_{l+1}^{t}|h_{l}^{0},h_{l}^{1},h_{l}^{2},\dots,h_{l}^{t}}$, simply denoted by $p_{h_{l+1}^{t}|H_{l}^{\leqslant t}}$.
Hypothesis 1 omits normalization layers, MLP layers, and residual structures between transformer blocks, and thus the output of $l$-th block is directly considered as the input to $l+1$-th block ($H_{l+1}$), as shown in Figure 1. This omission is supported by previous research [30] which shows that the flow of information in LLMs is largely concentrated within the attention layers of transformer architecture.
Accordingly, we could categorize token variables within each sequence into two distinct categories: microscopic (micro) variables and macroscopic (macro) variables. A micro variable refers to a token which is solely influenced by a single token as the input. For instance, $h^{0}$ satisfies $p_{h^{0}_{l+1}|h^{0}_{l}}$. Whereas macro variables aggregate information from all micro variables and thus encompass tokens which are influenced by all the tokens within the sequence as the input. An example could be $h^{T-1}_{l+1}$ which satisfies $p_{h^{T-1}_{l+1}|h^{0}_{l},h^{1}_{l},\dots,h^{T-1}_{l}}$. In summary, the NTP mechanism can be viewed as a behavior that increasingly coarsens from the most micro to the most macro scale.
With emergence described as “when quantitative changes in a system result in qualitative changes in behavior” [31], and the introduction of micro/macro variables, the definition within emergentism can be articulated as “Emergence is when micro changes in a fine-grained system result in macro changes in the collective perspective.” Succinctly, phenomenon observable at a macro level, yet unobservable at a micro level during a dynamic process, are abstracted as emergence.
@@
Given $T$ binary tokens $H_{l}=\{h^{0}_{l},h^{1}_{l},\dots,h^{T-1}_{l}\}\in\{0,1\}^{T}$ as inputs, for simplicity, we assume all variables are micro variables: $\forall h^{t}_{l+1}\in H_{l+1}$ satisfies $p_{h^{t}_{l+1}|h^{t}_{l}}$ (the simplest Markov process, and in the subsequent of this Example, we use $p$ to simply denote this transition probability). The output representations are also binary, i.e., $H_{l+1}\in\{0,1\}^{T}$. We assume an evolution rule which enables the parity of the sum of all output variables equal to the sum of all inputs with probability $\gamma$. If $H_{l}$ satisfies the uniform distribution, the evolution rule entails the probability of the output $H_{l+1}$ :
@@
With the assumption of micro dependency $p_{h^{t}_{l+1}|h^{t}_{l}}$, we can derive the distribution of a micro variable as $p(h^{t}_{l+1}$=$0|h^{t}_{l})=p(h^{t}_{l+1}$=$1|h^{t}_{l})=0.5$. Finally, let $h^{ma}$ be the macro variable with $h^{ma}$=$\uplus^{T-1}_{t=0}h^{t}$. Then the distribution of the macro variable becomes:
@@
Example 2 elucidates an interesting phenomenon known as emergence: The macro variable $h^{ma}$ is not induced by any individual micro variable but a collective of them. As a result, it shows different phenomenon from any micro variable. Moreover, we note that the same evolution rule applies to both micro and macro variables. This is known as Supervenience Hypothesis [4, 5] in emergentism:
Hypothesis 2 (Supervenience).
When the properties of the micro-level mechanisms of a system are fixed, so are the properties of all its macro levels.
Hypothesis 2 explicates that the “new phenomenon” in macro systems does not materialize ex nihilo. The micro level is causally and mechanically complete, and there is no room left for any causal and mechanical contribution at the macro level. In other words, the phenomenon that emerged in macro has irreducible causal and mechanical power in practice but not in principle, all because the micro level is too fine-grained to observe such a phenomenon.
Given the Supervenience hypothesis and motivated by more endeavors about emergentism [23, 13, 14], we can define emergence in LLMs as:
Definition 1 (Emergence in LLMs).
@@
For any transformer block $l$ in a LLM, let $h^{ma}_{l}$ be the macro variable ($h^{ma}_{l}$ satisfies $p_{h^{ma}_{l+1}|H^{\leqslant T}}$) and let $h^{mi}_{l}$ be the micro variable ($h^{mi}_{l}$ satisfies $p_{h^{mi}_{l+1}|h^{mi}_{l}}$), the strength of emergence in block $l$ can be described as:
$E(l)=MI(h^{ma}_{l+1},h^{ma}_{l})-\frac{1}{T}\sum^{T-1}_{t=0}MI(h^{mi\_t}_{l+1},h^{mi\_t}_{l})$ (3)
where $MI(\cdot,\cdot)$ represents the mutual information.
@@
Definition 1 describes how to estimate the strength of emergence. To illustrate, suppose an input sequence contains three tokens ‘large language model’, with their representations at the $l$th block denoted as $H_{l}=\{h^{0}_{l},h^{1}_{l},h^{2}_{l}\}$. To compute the mutual information at the micro level, we need to make sure each micro token is positioned at the beginning of the sequence to avoid the influence from other tokens due to the Markov process in Hypothesis 1. Specifically, we obtain the first micro variable $\{h^{mi\_0}_{l}\}^{L-1}_{l=0}$ where $h^{mi\_0}=h^{0}$ corresponds to ‘large’ derived by going through the transformer model for the input ‘large language model’. The second micro variable $\{h^{mi\_1}_{l}\}^{L-1}_{l=0}$ corresponds to ‘language’ given by going through the transformer model one more time for the modified input ‘language model’. The third micro variable proceeds in a similar manner by removing the first two tokens in the original input. These modified inputs ensure that each micro variable only depends on itself in the previous block. Meanwhile, the macro variable $\{h^{ma}_{l}\}^{L-1}_{l=0}$ is given by $h^{ma}_{l}=h^{2}_{l}$ for the original input sequence ‘large language model’. Finally, we have $E(l)=MI(h^{ma}_{l+1},h^{ma}_{l})-\frac{1}{3}(MI(h^{mi\_0}_{l+1},h^{mi\_0}_{l})+MI(h^{mi\_1}_{l+1},h^{mi\_1}_{l})+MI(h^{mi\_2}_{l+1},h^{mi\_2}_{l}))$.
Notably, We measure the emergence by comparing the macro changes with the micro changes across blocks, instead of directly computing $MI(h^{ma}_{l},h^{mi}_{l})$ at the same block, justified by the Supervenience hypothesis which posits that under the same state, micro and macro variables do not exhibit emergent properties. We use Example 2 to explain why the mutual information between macro variables and micro variables is different:
Example 2.
Continuing from Example 2, for simplicity, we assume $\gamma$=$1$, $T$=$3$. Assuming block $l$ undergoes a uniform distribution, we have $p(h^{ma}_{l}$=$0)=p(h^{ma}_{l}$=1)=0.51)=0.51 ) = 0.5 and $p(h^{mi\_t}_{l}$=$0)=p(h^{mi\_t}_{l}$=1)=0.51)=0.51 ) = 0.5. Without loss of generality, let’s assume $h^{ma}_{l}=0$ and $h^{mi\_t}_{l}=0$. According to Example 2, we have $p(h^{mi\_t}_{l+1}$=$0|h^{mi\_t}_{l})=p(h^{mi\_t}_{l+1}$=$1|h^{mi\_t}_{l})=0.5$. On the contrary, $p(h^{ma}_{l+1}=0|h^{ma}_{l})=\gamma=1$. Hence, the emergence value becomes $E(l)=1-\frac{1}{3}(0+0+0)=1$.
@@
Example 2 elucidates that the difference in mutual information stems from a coarse-graining that transitions from individual elements (micro) to an aggregate whole (macro). From an information theory perspective, $E(l)>0$ indicates that when the function of transformer block $l$ results in a higher reduction of uncertainty (entropy) on the whole sequence (macro variable) compared to the individual tokens (micro variables), there is a higher chance of observing an emergence phenomenon.
Consequently, the strength of emergence can be briefly understood as “how confidence with which a language model, based on previous tokens, definitely predicts the next token with a lower entropy in semantics”.
3.2 How to Estimate Emergence in LLMs?

### v2

# Quantifying Semantic Emergence in Language Models

@@ Abstract
Large language models (LLMs) are widely recognized for their exceptional capacity to capture semantics meaning. Yet, there remains no established metric to quantify this capability. In this work, we introduce a quantitative metric, Information Emergence (IE), designed to measure LLMs’ ability to extract semantics from input tokens. We formalize “semantics” as the meaningful information abstracted from a sequence of tokens and quantify this by comparing the entropy reduction observed for a sequence of tokens (macro-level) and individual tokens (micro-level). To achieve this, we design a lightweight estimator to compute the mutual information at each transformer layer, which is agnostic to different tasks and language model architectures. We apply IE in both synthetic in-context learning (ICL) scenarios and natural sentence contexts. Experiments demonstrate informativeness and patterns about semantics. While some of these patterns confirm the conventional prior linguistic knowledge, the rest are relatively unexpected, which may provide new insights. Our codes are available at: https://github.com/Zodiark-ch/Emergence-of-LLMs.
@@ 1 Introduction
One of the most elusive and captivating attributes of large language models (LLMs) is their ability to learn semantics from inputs across diverse domains [9, 7, 18]. However, it is unclear how to quantitatively measure the capability of LLMs in capturing semantics from texts.
Numerous existing tasks indirectly reflect similar capabilities through evaluating LLMs’ performances (e.g., accuracy) on a specific task, such as “instruction following” [30], “searching” [24], and “reasoning” [28]. Nevertheless, these evaluation methods rely on manually curating datasets and tasks tailoring different aspects, resulting in time-consuming and domain-specific findings. In addition, these evaluations typically focus on coarse-grained text, not providing interpretations for the behavior of finer-grained tokens. Lastly, existing evaluation metrics which vary across different tasks can lead to varied performances, and even contradicting conclusions [21].
In response to the above limitations, we propose a task-agnostic and closed-form metric, which we refer to as Information Emergence (IE) [^1] The “Information Emergence” here and “Emergence” in LLM-related research are two different notions, we discuss their difference in Section 2.2 and Section 5.3., designed to reflect and deterministically quantify the ability of LLMs to extract meaningful semantics from input tokens. To begin with, we construct a mathematical formalism capable of modeling semantics. In essence, semantics naturally emerge as a meaningfully organized ensemble of tokens [11, 3]. Consequently, tokens are considered microscopic (micro) observations with sophisticated patterns in a sentence, whereas semantics represent macroscopic (macro) observations emerging with more predictable behaviors. Inspired by information theory [4, 5], we formalize the model’s proficiency in semantics understanding, i.e., information emergence, as the difference of the entropy reduction between micro-level and macro-level. In another word, a better model proficient in deriving semantics from tokens, in comparison to other models, ought to render a higher entropy reduction for a global sequence than for a single token.
To compute IE in transformer models, as discussed earlier, we need to mathematically measure entropy reduction for both micro and macro levels. Given the auto-regressive nature of the next-token-prediction (NTP) mechanism, at any layer $l$ in transformer, the most micro-level transition can be naturally framed as the probability $p_{h^{0}_{l}|h^{0}_{l-1}}$ for an isolated token $h^{0}_{l}$, whereas the macro-level transition can be formulated as $p_{h^{T}_{l}|h^{0}_{l-1},h^{1}_{l-1},...,h^{T}_{l-1}}$ across $T$ tokens. We resort to the mutual information between successive transformer layers and adopt a practically effective estimation algorithm motivated by [6]. Therefore, we can measure the IE value for any token at any transformer layer, reflecting the strength of the LLM’s capability in extracting semantics from the historical context.
To validate the effectiveness of IE, we devise a suite of comprehensive experiments encompassing two different scenarios. In the first scenario, we curate a group of synthetic datasets under the ICL setting with different context domains. In the second scenario, we collect two wild datasets consisting of real-world natural language questions/answers. Under both scenarios, we experiment with different LMs including GPT-2 [19], GEMMA [25], and OpenLlama [10]. In alignment with our hypothesis, we show that IE offers a high-level informativeness through semantics faithfulness and sensitivity - the richer the semantics, the higher the IE. Furthermore, we obtain 3 interesting findings: 1) IE increases token-by-token in natural texts, whereas, in ICL-style texts, IE increases only when a new demonstration appears. 2) There is a strong correlation between specific hallucination phenomenons and a high variance in IE scores. 3) Distinctive patterns in IE have been observed between human-written and LLM-generated texts, revealing IE’s potential in automatically recognising LLM generations.
Overall, the main contributions could be summarized below:
  • We introduce IE, a novel, reasonably validated, and task-agnostic metric to deterministically quantify the semantic understanding capability of LLMs.
  • We introduce a light-weight implementation method for evaluating IE, which can be applied to extremely large and closed-source LMs like GPT-4 [1].
  • Empirical evidence demonstrates that IE can uncover previously unknown and essential patterns in areas such as ICL, Emergence, and hallucination.
@@ 2 Related Work
2.1 Evaluation on LLM Capabilities
The prevalent body of research extensively measures the capabilities of LLMs across various tasks by employing substantial benchmark datasets [23, 27, 33]. Additionally, a significant amount of research focuses on the performance of LLMs concerning specific capabilities such as adaptability to different domains [2], human-like cognition (opinions, attitudes, etc.) [17], followed with input instructions [30], text searching capability [24], and reasoning ability [28]. In contrast to these studies, our work concentrates on an essential yet abstract ability of LLMs - the ability to extract semantics from tokens.
2.2 Information Emergence and Emergence
“Information Emergence" and “Emergence" are two concepts with similar names but entirely different meanings. Emergence is defined as a capability that does not exist in smaller models but appears in larger ones [23, 16, 29, 15]. Most commonly, as the model size increases, the performance on many tasks rapidly improves. IE is a concept defined and validated in Information Theory [4, 5]. It describes phenomena observable at the macroscopic level but unobservable at the microscopic level. Nevertheless, only limited experiments are conducted to reflect a similar pattern of abrupt improvement between IE and Emergence with increased model size.
@@ 3 Method
3.1 How to Model Semantics in LLMs
In this paper, we identify the transformer block as the fundamental unit for LMs [^2] We focus on decoder-only language models.. Specifically, we employ $l=0,1,\dots,L-1$ to index transformer blocks within a language model, where $L$ represents the total number of blocks. For instance, GPT-2 XL (1.6B parameters) comprises 12 blocks ($L=12$), and Gemma-2B totals 18 blocks ($L=18$). For any transformer block $l$, given an input sequence of token length $T$ and hidden state dimension $D$, the input representation is given by $H_{l}=\{h_{l}^{0},h_{l}^{1},h_{l}^{2},\dots,h_{l}^{T-1}\}$ and the output representation is $H_{l+1}=\{h_{l+1}^{0},h_{l+1}^{1},h_{l+1}^{2},\dots,h_{l+1}^{T-1}\}$, where $H\in\mathbb{R}^{T\times D}$ and $h^{t}\in\mathbb{R}^{1\times D}$. Without loss of generality, we hypothesize that the multi-layer blocks constitutes a Markov process.
Figure 1: The analogy of auto-regressive process in NTP to Markov process. Taking the output representation of token2222 in Block 00 ($h^{2}_{1}$) as an example, which receives information from input representations of $h^{0}_{0}$, $h^{1}_{0}$, and $h^{2}_{0}$,satisfying $p_{h^{2}_{l+1}|H_{l}^{\leqslant t}}=p_{h^{2}_{l+1}|h^{0}_{l},h^{1}_{l},h^{2}_{l}}$.
@@ 3.1 How to Model Semantics in LLMs
The auto-regressive process of NTP mechanism in multi-layer blocks undergoes a Markovian stochastic process following a transition probability of any $h^{t}_{l+1}$ with $p_{h_{l+1}^{t}|h_{l}^{0},h_{l}^{1},h_{l}^{2},\dots,h_{l}^{t}}$, simply denoted by $p_{h_{l+1}^{t}|H_{l}^{\leqslant t}}$.
To simplify the analogy for easier understanding, Figure 1 omits normalization layers, MLP layers, and residual structures between transformer blocks, and thus the output of $l$-th block is directly considered as the input to $l+1$-th block ($H_{l+1}$). However, in all our real implementations, we retain the exact transformer output at every layer, i.e., $h_{l+1}=h_{l}+attention(h_{l})+MLP(h_{l})$.
Accordingly, we could categorize token variables within each sequence into two distinct categories: microscopic (micro) variables and macroscopic (macro) variables. A micro variable refers to a token which is solely influenced by a single token as the input. For instance, $h^{0}$ satisfies $p_{h^{0}_{l+1}|h^{0}_{l}}$. Whereas macro variables aggregate information from all micro variables and thus encompass tokens which are influenced by all the tokens within the sequence as the input. An example could be $h^{T-1}_{l+1}$ which satisfies $p_{h^{T-1}_{l+1}|h^{0}_{l},h^{1}_{l},\dots,h^{T-1}_{l}}$.
In summary, the NTP mechanism can be viewed as a behavior that increasingly coarsens from the most micro to the most macro scale and finally forms meaningful semantics. Hence, the macro level represents the semantics level and the micro level indicates the token level. Our defined IE commences with the discovery that the transmission probabilities of macro and micro variables differ in dynamic processes. (instantiated in Example 1).
@@
Given $T$ binary tokens $H_{l}=\{h^{0}_{l},h^{1}_{l},\dots,h^{T-1}_{l}\}\in\{0,1\}^{T}$ as inputs, for simplicity, we assume all variables are micro variables: $\forall h^{t}_{l+1}\in H_{l+1}$ satisfies $p_{h^{t}_{l+1}|h^{t}_{l}}$ (the simplest Markov process, and in the subsequent part of this Example, we use $p$ to simply denote this transition probability). The output representations are also binary, i.e., $H_{l+1}\in\{0,1\}^{T}$. We assume an evolution rule which enables the parity of the sum of all output variables equal to the sum of all inputs with probability $\gamma$. If $H_{l}$ satisfies the uniform distribution, the evolution rule entails the probability of the output $H_{l+1}$ :
@@
With the assumption of micro dependency $p_{h^{t}_{l+1}|h^{t}_{l}}$, we can derive the transition probability of a micro variable as $p(h^{t}_{l+1}$=$0|h^{t}_{l})=p(h^{t}_{l+1}$=$1|h^{t}_{l})=0.5$. Finally, let $h^{ma}$ be the macro variable with $h^{ma}$=$\uplus^{T-1}_{t=0}h^{t}$. Then the transition probability of a macro variable becomes:
@@
Example 1 elucidates an interesting phenomenon of IE: The macro variable $h^{ma}$ is not induced by any individual micro variable but a collective of them. As a result, it shows different phenomenon from any micro variable [^3] In fact, researchers in dynamics have proposed the Supervenience hypothesis [4, 5] to support this conclusion, but this is beyond the scope of our discussion..
Furthermore, to quantify the difference in transmission probabilities at the macro (semantics level) and micro (token level) in dynamic processes, we view it as a process of entropy reduction [20, 12, 13] and employ mutual information for modeling:
Definition 1 (Information Emergence in LLMs).
@@
For any transformer block $l$, let $h^{ma}_{l}$ be the macro variable ($h^{ma}_{l}$ satisfies $p_{h^{ma}_{l+1}|H_{l}^{\leqslant T}}$) and let $h^{mi}_{l}$ be the micro variable ($h^{mi}_{l}$ satisfies $p_{h^{mi}_{l+1}|h^{mi}_{l}}$), $MI(\cdot,\cdot)$ represents the mutual information, thus the strength of IE in block $l$ can be described as:
$E(l)=MI(h^{ma}_{l+1},h^{ma}_{l})-\frac{\sum^{T-1}_{t=0}MI(h^{mi\_t}_{l+1},h^{mi\_t}_{l})}{T}$ (3)
@@
Definition 3 describes how to estimate IE. To illustrate, suppose an input sequence contains three tokens ‘large language model’, with their representations at the $l$th block denoted as $H_{l}=\{h^{0}_{l},h^{1}_{l},h^{2}_{l}\}$. To compute the mutual information at the micro level, we need to make sure each micro token is positioned at the beginning of the sequence to avoid the influence from other tokens due to the auto-regressive nature of the NTP mechanism. Specifically, there are three micro variables ($\{h^{mi\_0}_{l}\}^{L-1}_{l=0}$, $\{h^{mi\_1}_{l}\}^{L-1}_{l=0}$, $\{h^{mi\_2}_{l}\}^{L-1}_{l=0}$) that correspond to ‘large’, ‘language’, and ‘model’ respectively. When calculating $\{h^{mi\_0}_{l}\}^{L-1}_{l=0}$, we treat a single token as the input text (that is, the input is ‘large’). Similarly, the other two micro variables are computed when ‘language’ and ‘model’ are respectively treated as the input text. These modified inputs ensure that each micro variable only depends on itself in the previous block. Meanwhile, the macro variable $\{h^{ma}_{l}\}^{L-1}_{l=0}$ is given by $h^{ma}_{l}=h^{2}_{l}$ for the original input sequence ‘large language model’. Finally, we have $E(l)=MI(h^{ma}_{l+1},h^{ma}_{l})-\frac{1}{3}(MI(h^{mi\_0}_{l+1},h^{mi\_0}_{l})+MI(h^{mi\_1}_{l+1},h^{mi\_1}_{l})+MI(h^{mi\_2}_{l+1},h^{mi\_2}_{l}))$.
@@
From an information theory perspective, $E(l)>0$ indicates that when the function of transformer block $l$ results in a higher reduction of uncertainty (entropy) on the whole sequence (macro variable) compared to the individual tokens (micro variables), there is a higher chance of capturing the collective semantics. Consequently, IE can be briefly understood as “how confident with which a language model, based on previous tokens, definitely predicts the next token with a lower entropy in semantics”.
3.2 How to Estimate IE in LLMs?